### v1

# LayerCollapse: Adaptive compression of neural networks

Abstract
Handling the ever-increasing scale of contemporary deep learning and transformer-based models poses a significant challenge. Although great strides have been made in optimizing model compression techniques such as model architecture search and knowledge distillation, the availability of data and computational resources remains a considerable hurdle for these optimizations. This paper introduces LayerCollapse, a novel alternative adaptive model compression methodology. LayerCollapse works by eliminating non-linearities within the network and collapsing two consecutive fully connected layers into a single linear transformation. This approach simultaneously reduces both the number of layers and the parameter count, thereby enhancing model efficiency. We also introduce a compression aware regularizer, which compresses the model in alignment with the dataset quality and model expressiveness, consequently reducing overfitting across tasks. Our results demonstrate LayerCollapse’s effective compression and regularization capabilities in multiple fine-grained classification benchmarks, achieving up to 74% post training compression with minimal accuracy loss. We compare this method with knowledge distillation on the same target network, showcasing a five-fold increase in computational efficiency and 8% improvement in overall accuracy on the ImageNet dataset.
1 Introduction
The expanding scope of Deep Learning (DL) applications has led to an escalating demand for more accurate models, resulting in an ongoing rise in their complexity. This surge in complexity, however, presents a challenge: it increases the execution costs, making it difficult to deploy these models in real-world scenarios on standard hardware. For example, transformer-based deep models while improving accuracy, present new computational challenges. Fortunately, modern trained neural networks are shown to include excess redundancies that can be removed without losing accuracy. This potential for reduction has catalyzed research in model compression strategies such as pruning [24][19], quantization [15][18][29][20], and model distillation[14], focusing on the effective detection and elimination of these superfluous components. Despite the amazing progress in the field of model compression, these optimizations demand the availability of data and high computational resources which further hinders their practicability.
Multi-Layer Perceptrons (MLPs) [11] is a key architectural component of modern deep neural networks. MLPs are regarded by many as the first contribution to deep learning, first introduced in 1965 by Ivakhnenko [17][32]. In its simplest form, a $K$-layer MLP can be described as $K$ linear transformations separated by some non-linear operations. Equivalently, an MLP can also be described as a sequence of fully connected (a.k.a., dense), layers separated by non-linear layers. Often, MLP layers serve as the final decision-making components, such as classification heads. MLPs are central to the success of deep learning models in diverse architectures including Convolutional Neural Networks [33], Vision Transformers [7], and MLP-Mixers [35].
Because of the dense nature of MLPs, a significant portion of a model’s computational overhead and storage requirements can be attributed to the MLP layers. For instance, MLPs constitute over 60% of the model’s total parameters for the popular models discussed in [7][35][33].
Table 1: Percentage of model parameters and MACs reductions attainable by collapsing their MLP layers
Model % Params % MACs
ViT T/16 $54.3$ $56.7$
ViT B/16 $57.3$ $57.9$
ViT L/16 $57.9$ $58.2$
Mixer B/16 $85.0$ $85.2$
Mixer L/16 $86.5$ $86.5$
VGG 16 $71.2$ $0.6$
MLPs contain non-linear activations, e.g., ReLU layers, which separate sequences of linear transformations. If the non-linear layers were absent, the consecutive linear layers could be condensed into a single layer.To showcase the immense potential of computation and storage reductions when linearizing MLPs, we calculate this reduction in the architectures presented in [7][35][33]. Table 1 demonstrates the reduction percentage in the number of parameters and MAC operations when we replace the MLP’s with a linear transform in relevant vision architectures.
The core thesis of our work revolves around non-linear activations. We argue that the expressiveness of an MLP is primarily governed by its non-linear components. By modulating these non-linear characteristics, we can orchestrate a continuum of model expressiveness, ranging from highly complex to near-linear transformations. Figure 1 demonstrates this idea using a simple regression task and an MLP with 2 layers. As illustrated, having complex models with lack of a sizeable dataset will lead to overfitting. By modulating the non-linear activation layer one can regulate model’s expressiveness.
Figure 1: Illustration of regression performance on a two-layer neural network, demonstrating the impact of varying levels of ReLU activation linearity on model overfitting and underfitting.
Classically, regularization strategies such as the $\ell_{1}$ and $\ell_{2}$ norms, have been employed to restrict model expressiveness, ensuring robust generalization and preventing overfitting. These techniques also contribute to compression-aware training. In similar veins, Lemaire et. al. describe a budget aware regularizer [23] and Nguyen et. al. introduce a regularizer for quantization aware training [29]. Extending this principle, we introduce a compression aware regularizer by using the nature of ReLU activations. This regularizer works by shifting the slope of Relu towards linearity. This not only facilitates the fusion of successive linear layers, conserving computational resources and storage, but also can lead to better model generalizability.
This paper presents several key contributions to the field of neural network compression and regularization:
  • Introduction of a novel post-training technique, called LayerCollapse, that combines two fully connected layers into one, reducing model size and complexity.
  • Development of a novel compression aware regularization loss with the objective of promoting shallower structures during training.
  • Proof-of-concept experimentation on a diverse set of datasets benchmarks, in vision and NLP domaains, showing post training compression up to 74%. Furthermore we show improvement in generalizability when training with proposed compression aware regularizer while decreasing size upto 50%. Lastly we compare the computational performance of knowledge distillation and LayerCollapse showing an 8% final result improvement while using 80% less computational resources.
  • Providing an open-source library that enables easy implementation of LayerCollapse, compatible with PyTorch. [^1] https://github.com/sohail2810/LayerCollapse
2 Related work
2.1 Compression
In the realm of deep learning, compression techniques aim to reduce computational demands while ensuring model precision. With the emergence of foundational models and the tapering of Moore’s law, the correct utilization of resources is a paramount concern. Predominantly, the methods employed for model compression are pruning [10], quantization [15][18], and distillation [14], each with its unique approach toward model optimization.
Pruning eliminates neurons whose impact on model performance are marginal or negligible. A big body of work explores diverse strategies to discern which neurons are expendable, including magnitude [24] and gradient [26][41] based methods. Another body of work looks into pruning aware training (PAT). In these methods a regularization term ensures structural sparsity when training [23][19][38][1]. This reduces the loss of accuracy when pruning the model post training.
Quantization constrains the parameter range and maps it into integers, thereby reducing the model’s computational and storage requirements. This methodology has been explored and implemented in various studies, providing numerous approaches toward effective quantization [9][29]. To minimize the loss induced by quantization it is standard practice to do quantization aware training (QAT). In short, we regularize the model weights during training to stay within a a close range of their quantized value [29].
Distillation[14], within another compression category, initially seeks an optimized architecture via Neural Architecture Search (NAS), a procedure notably resource-intensive in itself. The target model is trained using this new architecture, utilizing the dataset and the original model as a teacher. This process is sometimes called the student-teacher paradigm. It’s important to note that both NAS and distillation are resource-draining processes and require access to training data, which may not always be accessible.
In section 4, we compare LayerCollapse against distillation, highlighting its respective merits and applicability.
2.2 Regularization
In the field of deep learning, achieving the right balance between underfitting and overfitting is a pivotal challenge. Overfitting occurs when a model excessively learns from the noise in the training data, reducing its ability to generalize to new datasets. This issue is typically managed through the bias-variance trade-off, a principle that aids in creating models capable of effective generalization. A model with high variance is overly sensitive to small fluctuations in data, leading to unstable predictions. This trade-off highlights the necessity of maintaining an equilibrium between model complexity and the volume and integrity of the data. To counteract overfitting and enhance performance on test data, it’s crucial to implement explicit regularization strategies. These methods ensure the model’s robustness and consistent performance across both training and test datasets. Research by Yang et al. demonstrated that $\ell_{1}$ regularization can serve as a pruning-aware training method to sparsify the model, thereby boosting its pruning efficiency[40]. Ding et al. employed a specialized $\ell_{2}$ regularization to selectively eliminate less significant channels during training, aiming for subsequent pruning [4]. Additionally, Nguyen et al. introduced a regularization technique specifically for quantization-aware training [29]. Drawing inspiration from Santos et al.’s classification framework, we categorize model generalization techniques into three distinct categories: “Ensembling”, “Explicit Model Regularizers”, and “Dataset Design”[31].
Ensembling techniques often revolve around the integration of multiple models or sub-models to enhance performance and generalization. Within this domain, methods such as Dropout and its variants [34][5][8] play a pivotal role.
Dataset Design centers on refining the input data and labels to achieve model generalization. This includes data augmentation techniques such as transformative actions (random cropping, blurring, and rotation) seen in works like Bag of Tricks [13], FixRes [36], Cutout [3], CutMix [42], and AutoAugment [25].
Explicit regularizers modulate model configurations either through architectural adjustments or by penalizing certain weights to prevent overfitting. These methods follow the principle of Occam’s razor. They assume a prior knowledge that simpler models are more likely to be robust. This category encompasses classical regularization techniques like L0, L1, and L2 (weight decay) [27]. Our proposed method, falls into the same category. In section 3, we demonstrate how we use this principle in designing our regularization term. While classical norms influence weight magnitudes to mitigate overfitting, LayerCollapse uniquely stands out in its ability to adjust the model’s depth and structure.
3 Methodology
Figure 2: Overview of the LayerCollapse process: $(a)$ The compression aware the regularization loss shifts the PReLU towards linearity. $(b)$ The layer collapse operation involves eliminating a layer by replacing the hidden layer and adjusting the weights through the product of two matrices.
This section introduces LayerCollapse, an innovative compression method that adaptively reduces the model’s depth through regularization. This methodology aims to optimize the size of the model based on the quality of the dataset, resulting in enhanced generalization. Our proposed framework seamlessly integrates regularization and compression.
Unlike pruning [24] and quantization [15] which require the compressed model to be finetuned, LayerCollapse requires the finetuning prior to the compression phase.
As depicted in Figure 2, during a compression aware training or finetuning, we regularize a layer’s complexity by linearizing its activation function. We define Collapsing an MLP on a model that has shifted at least one of its non-linear layers into linearity. As depicted in Figure 2 Linearizing a non-linear activation within a model allows for the collapse of the two surrounding linear layers into one linear layer, effectively reducing the depth. The resulting model has fewer parameters and a smaller depth, requiring less resources for execution. For post training compression, one needs to run a few cycles of finetuning using the LayerCollapse regularizer to ensure the linearization loss is marginal.
We define a Collapsible MLP to be two consecutive fully connected layers utilizing a Parametric Rectified Linear Unit (PReLU) [12] in the middle. The Parametric Rectified Linear Unit (PReLU) can be expressed as:
$PReLU_{\alpha}(x)=\max(0,x)+\alpha\min(0,x)$ (1)
where $\alpha$ represents the slope of the PReLU function when x is negative. Output of a basic two-layer collapsible MLP is described as:
$Y_{\alpha}=W_{2}(PReLU_{\alpha}(W_{1}X+b_{1}))+b_{2}$ (2)
where, $W_{1}$ and $W_{2}$ are the weight matrices and $b_{1}$ and $b_{2}$ are the bias vectors for the first and second fully connected layers, respectively. $\alpha$ can be used to make the activation linear. We define $Y_{linear}$ as:
$Y_{linear}=W_{2}W_{1}X+W_{2}b_{1}+b_{2}$ (3)
We define the error as the squared difference of $Y_{\alpha}$ and $Y_{linear}$. In Appendix A, we show that $\forall\delta>0$
$P\left\{|Y_{\text{linear}}-Y_{\alpha}|^{2}\leq C\times(1-\alpha)^{2}\right\}>1-\delta$ (4)
where,
$x^{\delta}\equiv\inf\left\{x\in\mathcal{R}^{n}\mid P\left\{|X|>|x|\right\}<\delta\right\}$ (5)
$C=\sigma_{max}(W_{2}W_{1})^{2}|x^{\delta}|^{2}+|W_{2}b_{1}|^{2}$
where, $\sigma_{max}(W_{2}W_{1})$ is the largest singular value of $W_{2}W_{1}$ By choosing $\delta$ arbitrarily small, error is bounded proportional to $(1-\alpha)^{2}$ with probability one. We capitalize on this by incorporating $(1-\alpha)^{2}$ into our loss function. This loss term serves to align $Y_{\alpha}$ with $Y_{linear}$, compressing the model into a shallower model as demonstrated in Figure 2. In simpler terms, when $\alpha$ approaches one, substituting $Y_{linear}$ with $Y_{\alpha}$ with an error bounded by a constant multiplied by $(1-\alpha)^{2}$.
In practical applications, Dropout [34] and BatchNormalization [16] are frequently used between the layers of a network. Dropout works by temporarily deactivating random neurons during training, which helps spread the learning across different neurons to avoid over-reliance on any particular subset of them. It is important to note that Dropout is not active during model deployment, meaning it behaves as if it is not there. This characteristic means that Dropout doesn’t affect the layer collapsing process and can be left out when considering it.
BatchNormalization is a process described by the formula:
$Y=\frac{x-E[X]}{\sqrt{Var[X]}}\times\gamma+\beta$ (6)
where, $\gamma$ and $\beta$ are learnable parameters, the $E[X]$ and $Var[X]$ are the mean and variance of the current batch, which we sometimes adjust with a method called Polyak averaging to keep the batch-to-batch changes smooth. Algorithm 1 outlines the specific steps to be taken for adjusting weights and biases when BatchNormalization is present.
To measure how well our compression method works, we use the concept of “compression gain”. This is calculated as the percentage reduction in the number of parameters of a layer, excluding bias and BatchNormalization parameters. The formula for compression gain is:
$Gain=1-\frac{n_{in}\times n_{out}}{n_{hidden}\times(n_{in}+n_{out})}$ (7)
where, $n_{in}$, $n_{hidden}$, and $n_{out}$ denote the input size, hidden size, and output size of a 2-layer MLP, respectively.
An interesting aspect of this formula is that the gain can sometimes be negative. Positive gain happens when $n_{hidden}$ is larger than the product of $n_{in}$ and $n_{out}$ divided by their sum. This means that MLPs designed with a “bottleneck” structure (where $n_{hidden}$ is relatively small) may not see much benefit from our layer collapsing method. On the other hand, most modern MLP designs, which tend to expand the layer size (“widening” structure), are likely to experience significant positive gains from our method.
3.1 Adaptation for Convolutional Layers
Convolutional layers play a pivotal role in numerous neural network architectures. Essentially convolution can be formulated as sparse fully connected layers with some redundancies. Their nature as linear transforms permits representation via circulant matrices. A circulant matrix is a matrix where the row vectors have the same elements and each row vector is rotated one element to the right relative to the preceding row vector. An interesting property of these matrices is that their product also forms a circulant matrix, resulting in another convolution layer. The kernel size of the resulting convolution layer, denoted as $k$, is calculated by $k=k_{1}+k_{2}-1$ where, $k_{1}$ and $k_{2}$ are the kernel sizes of the original convolution layers. The compression gain can be calculated by:
$Gain_{conv}=1-\frac{k_{1}^{2}\times c_{in}\times c_{hidden}+k_{2}^{2}\times c_{hidden}\times c_{out}}{(k_{1}+k_{2}-1)^{2}\times c_{in}\times c_{out}}$ (8)
where, $c_{in}$, $c_{hidden}$, and $k_{1}$ correspond to the input channels, output channels, and the kernel size of the initial convolution layer, respectively. Meanwhile, $c_{hidden}$, $c_{out}$ and $k_{2}$ are the parameters of the subsequent convolution layer. However, an essential note when applying LayerCollapse to common CNN architectures, such as ResNet, is that the gain often turns out to be negative. To achieve a positive gain, it is imperative for $c_{hidden}$ to be significantly larger than both $c_{in}$ and $c_{out}$. This observation highlights the need for careful configuration when attempting to collapse layers in such architectures. Overall, this framework offers a principled approach to explore the potential benefits of LayerCollapse on convolution layers, illuminating the optimization potential for convolutional neural networks.
3.2 Prior Distribution Regularization
We extend the use of the parametric ReLU (PReLU) function to develop a regularization method based on the maximum a posteriori (MAP) framework. Aligning with Occam’s razor, which favors simpler explanations, we propose to adopt a normal distribution with a mean of 1 as the prior distribution for $\alpha$. This setting effectively limits the two layers to a linear space, reducing their functional complexity, as illustrated in Figure 1. This choice is backed by empirical results given in section 4. MAP framework suggests adding negative log likelihood of the prior distribution of parameters to the loss function. This is referred to as the prior loss or regularization loss. Following this principal the added regularization term is:
$\mathcal{L}_{\text{reg}}=lc\times(1-\alpha)^{2},\text{}$ (9)
where $lc$ is the regularization constant.
Algorithm 1 Layer Collapse
1:function Collapse($W_{1},b_{1},W_{2},b_{2},\gamma,\beta,\bar{\mu},\bar{\sigma},\alpha,\tau$)
2:     if $\alpha\leq\tau$ then
3:         $s_{BN}\leftarrow\frac{\gamma}{\sqrt{\bar{\sigma}}}$
4:         $W_{\text{new}}\leftarrow s_{BN}\times W_{2}\cdot W_{1}$
5:         $b_{\text{new}}\leftarrow W_{2}\cdot(s_{BN}\times(b_{1}-\bar{\mu})+\beta)+b_{2}$
6:         return $W_{\text{new}},b_{\text{new}}$
7:     else
8:         return un-collapsible
9:     end if
10:end function
4 Experiments
In this section we evaluate our method on both post training compression, and compression aware training with our regularization. Initially, we explore the standalone capability of LayerCollapse in post training compression of streamlined vision models including vision transformer, mlp-mixer and VGG. We then compare LayerCollapse with Knowledge Distillation on the cost of training and achievable results on a target network. Subsequently, we demonstrate the effectiveness of compression aware training in an ablation study on mid-sized datasets: TinyImageNet [22], and CIFAR100 [21]. Further extending the scope of our investigation, we exhibit preliminary results on the BERT [2] language model to provide a proof-of-concept for the generalizability of LayerCollapse in domains beyond computer vision, such as natural language processing (NLP). These experiments aim to demonstrate LayerCollapse’s versatility and its potential to enhance efficiency and performance in various machine learning contexts.
Model Collapsed Layers No. of Params (M) Top1 Acc. (%) Top5 Acc. (%) MACs (G)
ViT T/16 0 5.72 75.35 93 1.08
1 5.46 75.79 93 1.02
2 5.20 75.35 92 0.97
3 4.94 74.42 92 0.91
ViT S/16 0 22.05 81.38 96 4.24
1 21.02 80.96 96 4.04
2 19.98 79.74 95 3.84
ViT B/16 0 86.57 81.43 96 16.85
1 82.44 82.17 96 16.04
2 78.3 80.31 95 15.23
ViT L/16 0 304.33 84.80 98 59.66
2 289.64 84.22 97 56.77
Mixer B/16 0 59.88 76.47 92 12.61
2 55.64 76.11 92 11.71
4 51.39 75.67 91 10.81
VGG19 0 143.67 74.21 92 19.65
2 45.11 71.22 88 19.55
VGG16 0 138.36 73.36 92 15.48
2 39.80 67.16 88 15.39
VGG13 0 133.05 71.58 90 11.32
2 34.49 67.21 88 11.22
VGG11 0 132.86 70.37 90 7.62
2 34.31 66.36 88 7.52
Table 2: Post training compression results, showcasing accuracy and compression attained via LayerCollapse across different layers in relevant architectures.
4.1 Compression
Our compression experiments are performed on ImageNet-1k classification task [30]. Unlike pruning [24] and quantization [15] compression methods which require finetuning on the compressed method to recover accuracy, LayerCollapse utilizes finetuning with the regularizer to ensure accuracy. Afterwards the compression phase will endure no accuracy loss. We first assess the stand-alone compression efficacy of LayerCollapse. This is followed by a comparative analysis of LayerCollapse and Knowledge Distillation, focusing on training efficiency and resource utilization.
4.1.1 Stand-Alone Compression
Dataset. We employ the ImageNet-1k [30] dataset for both fine-tuning and validation processes. It comprises over 1.2 million images for training and 50,000 for validation.
Implementation. We adapt pre-trained models from the TIMM library [39] for MLP-Mixer [35] and ViT [7], alongside VGG variants from torchvision [28]. ViT and Mixer architectures use GeLU activations in their original implementation. In target layers, these are converted to ReLU to align with the requirements of LayerCollapse. The fine-tuning follows a consistent procedure across models, using Stochastic Gradient Descent (SGD) with momentum and a learning rate of $5e^{-4}$ as the optimizer. The loss function is comprised of three parts, cross entropy between labels and result, KL divergence between the output and uncollapsed models output and the LayerCollapse regularization term (9) with $lc=0.2$. This process is capped at 20 epochs. Each layer for all ViT experiments use 5 epochs. Number of epochs for VGG architectures are given in Table 3. The MAC operations are all calculated by propagating a random $224\times 224$ image through the model. Further details are enumerated in Appendix B.
Sensitivity Analysis. To quantify the robustness of LayerCollapse, we conduct a layer-by-layer sensitivity analysis on the Vision Transformer T/16 model, as shown in Figure 3. This removal of layers provides insight into the impact of incremental compression on the model’s accuracy. Interestingly, we observe an initial increase in accuracy upon the collapse of the first layer, indicating an improvement in generalization compared to the original model.
Figure 3: Layer-wise collapse accuracy and parameter reduction analysis for ViT-T/16.
Model Method Top1 Acc(%) Top5 Acc(%) No. of Parameters (M) Cost Epochs
Initial Final
vgg19 LC 71.22 88 143.67 45.11 11h 5
KD 63.42 85 51h 23
vgg16 LC 70.01 89 138.36 39.80 14h 7
KD 59.63 82 31h 16
vgg13 LC 67.21 88 133.05 34.49 16h 10
KD 58.26 81 45h 28
vgg11 LC 66.36 88 132.86 34.31 20h 20
KD 58.64 81 42h 42
Table 3: A comparative analysis of Layer Collapse (LC) and Knowledge Distillation (KD) methods applied to VGG models.
Discussion. Table 2 illustrates that LayerCollapse achieves a substantial reduction in model size—up to 74% in VGG11—and computational demands with minimal accuracy degradation. Notably, models with heavier classification heads, such as VGG, can benefit significantly from our method. ViT models, the T/16 and B/16 variants, get an increase in accuracy with the first layer collapse. This can be attributed to an increase of generalization in simpler models. In ViT T/16, S/16, and B/16 we get roughly 5 percent reduction in model size per LayerCollapse. In all ViT variants we get a consistent reduction of over 10% in size and complexity with less than a 1% accuracy loss. In the Mixer architecture, 2 types of MLP layers known as MLP-token and MLP Channels exist. MLP-Channel has a significantly bigger size. We found that MLP-tokens can be collapsed in about an epoch of computation whilst MLP-channels require more effort. In the Mixer experiment, we collapsed the layers in pairs to ensure MLP-Channels are affected. In the Mixer architecture we get 14% reduction in size with less than a 1% accuracy degradation. Notice that the Mixer architecture is the only one that gets a linear reduction in MAC operations with each pair collapsed. We used the idea first introduced in Knowledge Distillation to enhance the performance of our finetuning procedure. Using the KL-Divergence loss, we speedup the gradient flow and increase the convergence speed.
4.2 LayerCollapse vs Knowledge Distillation
Setup. Our evaluation framework utilizes batch-normalized VGG variants from the torchvision repository [28], with ImageNet-1K [30] providing a substantial dataset for training and validation. These variants are chosen for their prevalent use in compression studies and their capability to learn without prior training. To facilitate a fair comparison between LayerCollapse and Knowledge Distillation, we adopt the same target architecture for both methods. The rest of the training is set up identical to section 4.1.1 for both LayerCollapse and Knowledge distillation.
Comparison. The models are rigorously compared across several dimensions: training cost, the number of training epochs required, and the accuracy metrics—top-1 and top-5—on the validation set. The computational costs are quantified on an Nvidia RTX A6000 GPU.
Discussion. Referencing Table 3, LayerCollapse consistently achieves over 10% higher accuracy while incurring only 30% of the computational cost compared to Knowledge Distillation. This trend is more pronounced in larger models, which exhibit greater compression potential, and result in up to 15% accuracy gain with 50% of tuning cost. It is noteworthy that for models necessitating pre-training, such as MLP-Mixer and ViT, the absence of pre-training data and the high computational demands often make Knowledge Distillation impractical. LayerCollapse emerges as a unique method for model compression through architectural modification, offering significant cost advantages as demonstrated in Table 3. However, it is important to recognize the inherent limitations of LayerCollapse in terms of the diversity of achievable target model architectures.
4.3 Ablation Study of Regularization
We provide an in-depth analysis of how our regularization approach impacts both the generalization and compression abilities in training image classifiers from scratch.
Dataset. Our study utilizes the CIFAR100 [10] and TinyImageNet [22] datasets, which are mid-sized and rich in diversity. The CIFAR100 dataset includes 60,000 images across 100 classes, and TinyImageNet, a subset of the larger ImageNet, comprises 100,000 images across 200 classes.
Implementation. The experiments are conducted using the VGG architecture because of its effective learning capabilities without pre-training and competitive performance in benchmarks. For consistency, all experiments utilize the same setup and hyper parameters. Note that the goal of this study is not to show state of the art performance using our proposed regularization, but to show its effectiveness in reducing over-fitting and final model size. In all experiments, traditional regularizers, such as data augmentation (rotation, random crop, and flipping), dropout[5], and batch normalization[16], have been used. This is to show the added contribution of 9. Training is done using Stochastic Gradient Descent (SGD) with momentum a learning rate scheduler, trained for a total of 400 epochs per experiment. The regularization is applied to 40% of the layers, starting with the latest layers, with the strength of 0.05 across all experiments. extensive experiment details can be found in Appendix B.
Discussion. The results, documented in Table 4, demonstrate the benefits of incorporating compression aware regularization. We observe an enhancement in top-1 accuracy by as much as 2% with 50% reduction in model size. The learning curve presented in Figure 4 suggests a narrowing gap between training and validation accuracy, indicating a reduction in overfitting and highlighting the effectiveness of our regularization strategy in improving generalization.
4.4 LayerCollapse in NLP
We extended LayerCollapse to the NLP domain. This section presents our initial findings using the BERT-Base model [2], a well-established pre-trained NLP model.
Setup In our evaluation we use the General Language Understanding Evaluation on the Microsoft Research Paraphrase Corpus (GLUE-MRPC)[37][6]. This benchmark assesses a model’s capacity to discern paraphrased sentences, determining if they are equivalent. Further details are found in Appendix B
Discussion We report our results in Table 5. These include accuracy metrics and the number of parameters for each layer of BERT-Base modified by LayerCollapse. The experiment shows a promising 19% reduction in parameters with a minimal accuracy loss of less than 3%. Notably, we exclusively use the data in the GLUE-MRPC task in the LayerCollapse procedure. These preliminary results in the NLP sector suggest the versatility and potential of LayerCollapse across different domains, not limited to computer vision. They point to the method’s adaptability and effectiveness in optimizing model parameters while preserving accuracy without the need for pre-train dataset.
Figure 4: Evaluating the impact of LayerCollapse regularization on CIFAR100, we compare the top-1 accuracy of VGG11 both with and without this regularization technique.
Table 4: Number of parameters after Collapse and Top-1 Accuracy of various models with and without LC on CIFAR-100 and Tiny ImageNet.
Model CIFAR-100 Tiny ImageNet
#Param Top-1(%) #Param Top-1(%)
VGG11 $11.7$M $59.7$ $18.4$M $60.2$
VGG11(LC) $9.3$M $63.5$ $9.6$M $61.7$
VGG13 $11.9$M $66.1$ $18.6$M $50.3$
VGG13(LC) $9.5$M $67.8$ $9.8$M $52.2$
VGG16 $17.2$M $63.3$ $23.9$M $48.9$
VGG16(LC) $14.8$M $64.1$ $15.1$M $49.7$
VGG19 $22.5$M $62.5$ $29.6$M $46.8$
VGG19(LC) $20.1$M $63.8$ $20.4$M $48.2$
Table 5: Layer-wise analysis of accuracy and parameters for Bert-Base via LC
Layers Accuracy (%) Params (M)
0 $85.0$ $109.40$
1 $85.8$ $105.35$
2 $84.1$ $101.20$
3 $84.3$ $97.10$
4 $82.1$ $92.95$
5 $82.3$ $88.80$
6 $82.3$ $84.70$
7 $80.4$ $80.56$
8 $79.4$ $76.40$
5 Conclusion
This paper introduces LayerCollapse, a novel method for compressing neural networks through regularization. LayerCollapse simplifies a network by merging pairs of linear layers, reducing the network’s depth, the number of parameters, thereby, reducing operational requirements. We introduce a compression aware regularizer that encourages the development of shallower network architectures.
Under rigorous evaluation, LayerCollapse demonstrates state-of-the-art performance in compressing Vision Transformers, improving even improving accuracy in some cases. In comparison to Knowledge Distillation, LayerCollapse consistently demonstrated superior efficiency in both computation and final results, achieving over 10% better performance while utilizing 80% fewer resources. LayerCollapse marks a significant paradigm shift in efficient neural network design and bridges the gap between performance and model size.
Supplementary Material
Appendix A Proof of Equation (4)
Given $Y_{linear}=W_{2}W_{1}X+W_{2}b_{1}+b_{2}$ and $Y_{\alpha}=W_{2}PReLU_{\alpha}(W_{1}X+b_{1})+b_{2}$ we show that $\forall\delta>0$
$P\left\{|Y_{\text{linear}}-Y_{\alpha}|^{2}\leq C\times(1-\alpha)^{2}\right\}>1-\delta$ (10)
where,
$x^{\delta}\equiv\inf\left\{x\in\mathcal{R}^{n}\mid P\left\{|X|>|x|\right\}<\delta\right\}$ (11)
$C=\sigma_{max}(W_{2}W_{1})^{2}|x^{\delta}|^{2}+|W_{2}b_{1}|^{2}$
where, $\sigma_{max}(W_{2}W_{1})$ is the largest singular value of $W_{2}W_{1}$.
We define indicator random variable $I$ as:
$I\equiv\begin{cases}1&W_{1}X+b_{1}\geq 0\\ 0&otherwise\end{cases}$ (12)
Writing $Y_{\alpha}$ using (12) we have:
$Y_{\alpha}=IY_{linear}+(1-I)(\alpha W_{2}W_{1}X+\alpha W_{2}b_{1}+b_{2})$ (13)
Using the definition of $Y_{linear}$, the squared distance of $Y_{linear}$ and $Y_{\alpha}$ can be written as:
$|Y_{\text{linear}}-Y_{\alpha}|^{2}=(1-I)^{2}(1-\alpha)^{2}(W_{2}W_{1}X+W_{2}b_{1})^{2}$ (14)
Since $(1-I)^{2}$ is either 00 or $1$ we have:
$|Y_{\text{linear}}-Y_{\alpha}|^{2}\leq(1-\alpha)^{2}(W_{2}W_{1}X+W_{2}b_{1})^{2}$ (15)
Using the triangle inequality we can write:
$|Y_{\text{linear}}-Y_{\alpha}|^{2}\leq(1-\alpha)^{2}(|W_{2}W_{1}X|^{2}+|W_{2}b_{1}|^{2})$ (16)
It is known that:
$|W_{2}W_{1}X|^{2}\leq\sigma(W_{2}W_{1})^{2}|X|^{2},$ (17)
where, $\sigma(W_{2}W_{1})$ is the largest singular value of the $W_{2}W_{1}$ matrix. Using (17) we have:
$|Y_{\text{linear}}-Y_{\alpha}|^{2}\leq(1-\alpha)^{2}(\sigma(W_{2}W_{1})^{2}|X|^{2}+|W_{2}b_{1}|^{2})$ (18)
Using $x^{\delta}$ and $\delta$ as defined in equation (11) we have:
[…]
$(1-\alpha)^{2}(\sigma(W_{2}W_{1})^{2}|x^{\delta}|^{2}+|W_{2}b_{1}|^{2})\}>1-\delta$
Using equations (19) and (18) with using the $C$ defined in (11) we have:
$P\left\{|Y_{\text{linear}}-Y_{\alpha}|^{2}\leq C\times(1-\alpha)^{2}\right\}>1-\delta$ (20)
Appendix B Experiment Details
The experimental files are accessible at [https://github.com/sohail2810/LayerCollapse].
B.1 Stand-Alone Compression
Dataset: We utilized the full ImageNet-1k dataset, with image dimensions of $224\times 224$. Training images were augmented with random cropping, horizontal flipping, and rotation (up to 15 degrees). Validation images were not augmented. Normalization factors were model-specific: standard and mean values were set to $0.5$ for all color channels in most models, except for the MLP-Mixer, which used the default ImageNet normalization (mean = ($0.482,0.456,0.406$), std = ($0.229,0.224,0.225$)). The batch size was fixed at 256 for all experiments.
Training Setup: We accelerated finetuning by using the original model’s predicted labels, akin to Knowledge Distillation techniques. The loss function combined cross-entropy and KL Divergence terms:
$loss=\frac{1}{2}CrossEntropy(p,\hat{p})+\frac{1}{2}$ $KLDivergance(p,\tilde{p})$ (21)
$+Regularization$
where $\hat{p}$ and $\tilde{p}$ represent true and original model’s predicted labels, respectively. Regularization is as defined in Eq. (9).
Layer collapsing was performed sequentially, starting from the last layer. Each layer’s MLP was replaced with our CollapsibleMLP, featuring PReLU activations, and finetuned using Eq. (21) for up to 10 epochs. Following this, we applied the collapse procedure (Algorithm 1). A regularization strength of $0.05$ and an SGD optimizer (momentum $0.9$, initial learning rate $0.0005$) were used. After 5 epochs, the learning rate was reduced to $10^{-4}$.
B.2 LayerCollapse vs. Knowledge Distillation
For both LayerCollapse and Knowledge Distillation, the target architectures were identical. The LayerCollapse setup followed the procedure outlined in Section B.1. The distillation process employed the SGD optimizer with an initial learning rate of $0.001$, reduced to $0.0001$ upon plateau. Experiments concluded when no further progress was achievable.
B.3 Ablation Study of Regularization
To ensure comparability, all models, including those without LayerCollapse, shared the same training settings. The SGD optimizer with a starting learning rate of $0.005$ was used, decreasing by a factor of $0.3$ at epochs 75, 150, and 250. A dropout probability of $0.5$ was applied. In LayerCollapse experiments, $0.4$ fraction of layers were regularized with a strength of $0.05$. Hyper-parameter tuning was not a focus of this study.
B.4 LayerCollapse in NLP
We utilized the base Bert model from Hugging Face for NLP experiments. Finetuning employed the Adam optimizer with a learning rate of $0.00002$.

### v3

@@
-LayerCollapse: Adaptive compression of neural networks
+
 Abstract
 Handling the ever-increasing scale of contemporary deep learning and transformer-based models poses a significant challenge.
-Although great strides have been made in optimizing model compression techniques such as model architecture search and knowledge distillation, the availability of data and computational resources remains a considerable hurdle for these optimizations.
-This paper introduces LayerCollapse, a novel alternative adaptive model compression methodology.
-LayerCollapse works by eliminating non-linearities within the network and collapsing two consecutive fully connected layers into a single linear transformation.
-This approach simultaneously reduces both the number of layers and the parameter count, thereby enhancing model efficiency.
-We also introduce a compression aware regularizer, which compresses the model in alignment with the dataset quality and model expressiveness, consequently reducing overfitting across tasks.
-Our results demonstrate LayerCollapse’s effective compression and regularization capabilities in multiple fine-grained classification benchmarks, achieving up to 74% post training compression with minimal accuracy loss.
-We compare this method with knowledge distillation on the same target network, showcasing a five-fold increase in computational efficiency and 8% improvement in overall accuracy on the ImageNet dataset.
+Overparameterized Transformer networks outperform prior art in Natural Language processing and Computer Vision.
+These models contain hundreds of millions of parameters, demanding significant computational resources and making them prone to overfitting on down stream tasks.
+In this work we present LayerCollapse , a novel structured pruning method to reduce the depth of fully connected layers.
+We propose an innovative regularizer that promotes shallow fully connected layers, compressing the model with minimal performance impact.
+This regularizer enables post-training compression without fine-tuning while preserving performance.
+LayerCollapse controls model expressiveness by regularizing the activation functions between fully connected layers, modulating them to linearity.
+A linear activation function collapses the rank of a transformation to the rank of the corresponding linear transformation, which demands less resources from the hardware.
+We demonstrate the effectiveness of LayerCollapse by showing its compression capabilities in sentimental analysis, text generation, and image classification benchmarks.
+footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
+Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
 1 Introduction
 The expanding scope of Deep Learning (DL) applications has led to an escalating demand for more accurate models, resulting in an ongoing rise in their complexity.
-This surge in complexity, however, presents a challenge: it increases the execution costs, making it difficult to deploy these models in real-world scenarios on standard hardware.
-For example, transformer-based deep models while improving accuracy, present new computational challenges.
-Fortunately, modern trained neural networks are shown to include excess redundancies that can be removed without losing accuracy.
-This potential for reduction has catalyzed research in model compression strategies such as pruning [24][19], quantization [15][18][29][20], and model distillation[14], focusing on the effective detection and elimination of these superfluous components.
-Despite the amazing progress in the field of model compression, these optimizations demand the availability of data and high computational resources which further hinders their practicability.
-Multi-Layer Perceptrons (MLPs) [11] is a key architectural component of modern deep neural networks.
-MLPs are regarded by many as the first contribution to deep learning, first introduced in 1965 by Ivakhnenko [17][32].
-In its simplest form, a $K$-layer MLP can be described as $K$ linear transformations separated by some non-linear operations.
-Equivalently, an MLP can also be described as a sequence of fully connected (a.k.a., dense), layers separated by non-linear layers.
-Often, MLP layers serve as the final decision-making components, such as classification heads.
-MLPs are central to the success of deep learning models in diverse architectures including Convolutional Neural Networks [33], Vision Transformers [7], and MLP-Mixers [35].
-Because of the dense nature of MLPs, a significant portion of a model’s computational overhead and storage requirements can be attributed to the MLP layers.
-For instance, MLPs constitute over 60% of the model’s total parameters for the popular models discussed in [7][35][33].
-Table 1: Percentage of model parameters and MACs reductions attainable by collapsing their MLP layers
+This surge in complexity, however, presents a challenge: It increases the execution costs, in terms of memory and latency.
+This makes it difficult to deploy these models in real-world scenarios on standard hardware.
+Recent studies have shown that modern neural networks often contain unnecessary redundancies, which can be eliminated without affecting their performance.
+This discovery has spurred interest in model compression techniques such as unstructured pruning Li et al.
+(2016) Jiang et al.
+(2023), structured pruning Fan et al.
+(2019), quantization Hubara et al.
+(2017) Jacob et al.
+(2018) Nguyen et al.
+(2020) Krishnamoorthi (2018), and model distillation Hinton et al.
+(2015).
+These methods aim to identify and remove the redundant elements effectively.
+However, despite significant advances in model compression, implementing the pertinent optimizations typically requires additional steps such as fine-tuning.
+This can be resource-intensive, making the compression techniques less practical.
+Moreover some of these methods like unstructured pruning require support from hardware and software to take advantage of the compression Sundaram et al.
+(2023).
+Table 1: Percentage of model parameters and MACs associated with MLP layers of streamline architectures.
 Model % Params % MACs
-ViT T/16 $54.3$ $56.7$
-ViT B/16 $57.3$ $57.9$
+GPT2-Large $53.4$ $52.7$
+Bert-Large $52.6$ $57.1$
 ViT L/16 $57.9$ $58.2$
-Mixer B/16 $85.0$ $85.2$
 Mixer L/16 $86.5$ $86.5$
 VGG 16 $71.2$ $0.6$
-MLPs contain non-linear activations, e.g., ReLU layers, which separate sequences of linear transformations.
-If the non-linear layers were absent, the consecutive linear layers could be condensed into a single layer.To showcase the immense potential of computation and storage reductions when linearizing MLPs, we calculate this reduction in the architectures presented in [7][35][33].
-Table 1 demonstrates the reduction percentage in the number of parameters and MAC operations when we replace the MLP’s with a linear transform in relevant vision architectures.
-The core thesis of our work revolves around non-linear activations.
-We argue that the expressiveness of an MLP is primarily governed by its non-linear components.
+Multi-Layer Perceptrons (MLPs) Haykin (1994) are a key architectural component of modern deep neural networks.
+MLPs are regarded by many as the first contribution to deep learning, present in modern deep learning designs such as the transformer architecture.
+In its simplest form, a $K$-layer is defined as as $K$ linear transformations separated by non-linear operations.
+MLPs are central to the success of diverse deep learning architectures including encoder only transformer models Vaswani et al.
+(2017) Lan et al.
+(2019) Raffel et al.
+(2020), causal transformer models Radford et al.
+(2019) Touvron et al.
+(2023) Gunasekar et al.
+(2023), Vision Transformers Dosovitskiy et al.
+(2020), and MLP-Mixers Tolstikhin et al.
+(2021).
+Due to the dense nature of MLPs, a significant portion of a model’s computational overhead and storage requirements can be attributed to its MLP layers.
+For instance, MLPs constitute over 60% of the model’s total parameters for the popular architectures discussed in Dosovitskiy et al.
+(2020) Tolstikhin et al.
+(2021) Simonyan & Zisserman (2014).
+MLPs contain non-linear activations, e.g., ReLU, which separate sequences of linear transformations.
+If the non-linear layers were absent, the consecutive linear layers could be condensed into a single linear transformation.
+Table 1 showcases the portion of storage and computation requirements of MLP layers is the mentioned architectures.
+Figure 1: Illustration of regression performance on a two-layer neural network, demonstrating the impact of varying levels of ReLU activation linearity on model overfitting and underfitting.
+We define the percent of linearity to be the negative slope of the ReLU activation.
+Figure 2: Overview of the LayerCollapse process: $(a)$ The regularization loss shifts the PReLU towards linearity.
+$(b)$ The layer collapse operation involves eliminating a layer by replacing the hidden layer and adjusting the weights through the product of two matrices.
+In this work, we focus on the non-linear activations inside MLPs.
+We argue that the expressiveness of an MLP is primarily governed by its non-linear component.
 By modulating these non-linear characteristics, we can orchestrate a continuum of model expressiveness, ranging from highly complex to near-linear transformations.
-Figure 1 demonstrates this idea using a simple regression task and an MLP with 2 layers.
-As illustrated, having complex models with lack of a sizeable dataset will lead to overfitting.
-By modulating the non-linear activation layer one can regulate model’s expressiveness.
-Figure 1: Illustration of regression performance on a two-layer neural network, demonstrating the impact of varying levels of ReLU activation linearity on model overfitting and underfitting.
-Classically, regularization strategies such as the $\ell_{1}$ and $\ell_{2}$ norms, have been employed to restrict model expressiveness, ensuring robust generalization and preventing overfitting.
-These techniques also contribute to compression-aware training.
-In similar veins, Lemaire et.
-al.
-describe a budget aware regularizer [23] and Nguyen et.
-al.
-introduce a regularizer for quantization aware training [29].
-Extending this principle, we introduce a compression aware regularizer by using the nature of ReLU activations.
-This regularizer works by shifting the slope of Relu towards linearity.
-This not only facilitates the fusion of successive linear layers, conserving computational resources and storage, but also can lead to better model generalizability.
-This paper presents several key contributions to the field of neural network compression and regularization:
-  • Introduction of a novel post-training technique, called LayerCollapse, that combines two fully connected layers into one, reducing model size and complexity.
-  • Development of a novel compression aware regularization loss with the objective of promoting shallower structures during training.
-  • Proof-of-concept experimentation on a diverse set of datasets benchmarks, in vision and NLP domaains, showing post training compression up to 74%.
-Furthermore we show improvement in generalizability when training with proposed compression aware regularizer while decreasing size upto 50%.
-Lastly we compare the computational performance of knowledge distillation and LayerCollapse showing an 8% final result improvement while using 80% less computational resources.
-  • Providing an open-source library that enables easy implementation of LayerCollapse, compatible with PyTorch. [^1] https://github.com/sohail2810/LayerCollapse
-2 Related work
-2.1 Compression
-In the realm of deep learning, compression techniques aim to reduce computational demands while ensuring model precision.
-With the emergence of foundational models and the tapering of Moore’s law, the correct utilization of resources is a paramount concern.
-Predominantly, the methods employed for model compression are pruning [10], quantization [15][18], and distillation [14], each with its unique approach toward model optimization.
-Pruning eliminates neurons whose impact on model performance are marginal or negligible.
-A big body of work explores diverse strategies to discern which neurons are expendable, including magnitude [24] and gradient [26][41] based methods.
-Another body of work looks into pruning aware training (PAT).
-In these methods a regularization term ensures structural sparsity when training [23][19][38][1].
-This reduces the loss of accuracy when pruning the model post training.
-Quantization constrains the parameter range and maps it into integers, thereby reducing the model’s computational and storage requirements.
-This methodology has been explored and implemented in various studies, providing numerous approaches toward effective quantization [9][29].
-To minimize the loss induced by quantization it is standard practice to do quantization aware training (QAT).
-In short, we regularize the model weights during training to stay within a a close range of their quantized value [29].
-Distillation[14], within another compression category, initially seeks an optimized architecture via Neural Architecture Search (NAS), a procedure notably resource-intensive in itself.
-The target model is trained using this new architecture, utilizing the dataset and the original model as a teacher.
-This process is sometimes called the student-teacher paradigm.
-It’s important to note that both NAS and distillation are resource-draining processes and require access to training data, which may not always be accessible.
-In section 4, we compare LayerCollapse against distillation, highlighting its respective merits and applicability.
-2.2 Regularization
-In the field of deep learning, achieving the right balance between underfitting and overfitting is a pivotal challenge.
-Overfitting occurs when a model excessively learns from the noise in the training data, reducing its ability to generalize to new datasets.
-This issue is typically managed through the bias-variance trade-off, a principle that aids in creating models capable of effective generalization.
-A model with high variance is overly sensitive to small fluctuations in data, leading to unstable predictions.
-This trade-off highlights the necessity of maintaining an equilibrium between model complexity and the volume and integrity of the data.
-To counteract overfitting and enhance performance on test data, it’s crucial to implement explicit regularization strategies.
-These methods ensure the model’s robustness and consistent performance across both training and test datasets.
-Research by Yang et al.
-demonstrated that $\ell_{1}$ regularization can serve as a pruning-aware training method to sparsify the model, thereby boosting its pruning efficiency[40].
-Ding et al.
-employed a specialized $\ell_{2}$ regularization to selectively eliminate less significant channels during training, aiming for subsequent pruning [4].
-Additionally, Nguyen et al.
-introduced a regularization technique specifically for quantization-aware training [29].
-Drawing inspiration from Santos et al.’s classification framework, we categorize model generalization techniques into three distinct categories: “Ensembling”, “Explicit Model Regularizers”, and “Dataset Design”[31].
-Ensembling techniques often revolve around the integration of multiple models or sub-models to enhance performance and generalization.
-Within this domain, methods such as Dropout and its variants [34][5][8] play a pivotal role.
-Dataset Design centers on refining the input data and labels to achieve model generalization.
-This includes data augmentation techniques such as transformative actions (random cropping, blurring, and rotation) seen in works like Bag of Tricks [13], FixRes [36], Cutout [3], CutMix [42], and AutoAugment [25].
-Explicit regularizers modulate model configurations either through architectural adjustments or by penalizing certain weights to prevent overfitting.
-These methods follow the principle of Occam’s razor.
-They assume a prior knowledge that simpler models are more likely to be robust.
-This category encompasses classical regularization techniques like L0, L1, and L2 (weight decay) [27].
-Our proposed method, falls into the same category.
-In section 3, we demonstrate how we use this principle in designing our regularization term.
-While classical norms influence weight magnitudes to mitigate overfitting, LayerCollapse uniquely stands out in its ability to adjust the model’s depth and structure.
+To gain intuition on the effect of non-linear activations control on model expressiveness, we use a simple 2-layer MLP for a simple regression task.
+The results are illustrated in figure 1.
+Here an $n\%$ linear model is defined as a model where the negative slope of the leaky ReLU is $n\%$ of one.
+So, the regular model is just a 2-layer MLP with ReLU activation and a $98\%$ linear model is a 2-layer MLP with a leaky ReLU activation with a negative slope of $0.98$.
+As we can see, different $n$ values create a continuum of model expressiveness.
+The regular model overfits to the noise within the data, however, the $80\%$ linear model fits the underlying distribution nicely while the $98\%$ linear model under-fits the data and cannot capture the complexity of the regression task.
+A linear transformation can be performed using only a single matrix multiplication, while a $K$-layer MLP requires $K$ matrix multiplications.
+By adjusting the non-linear component towards linearity, we can regulate the model expressiveness while reducing the operational complexity of the model.
+We replace two matrix multiplications and a non-linear transform with only one smaller matrix multiplication when the modulated non-linearity is close to linear.
+In section 3, we will formally define close to linear and its effect on accuracy.
+Controlling the model expressiveness using this technique will allow us to tailor the complexity of our model based on the task.
+Starting from the same architecture, some tasks have small datasets that require a less expressive model to prevent overfitting.
+In this work, we developed a novel regularization technique to limit the model expressiveness to prevent overfitting, while reducing the computation cost of the model.
+Common, regularization strategies such as Dropout Srivastava et al.
+(2014), $\ell_{1}$ and $\ell_{2}$ norms, have been employed to restrict model expressiveness, attempting robust generalization and preventing overfitting.
+These regularizers change the model weight structures such that post training compression like pruning and quantization pared with finetuning will yield efficient high performing models.
+Fan et al.
+introduced LayerDrop, a regularization technique to allow for post-training compression without fine-tuning Fan et al.
+(2019).
+LayerDrop employs a regularizer that controls the amount of bypass signals in each MLP layer within transformer architectures.
+Extending this idea, we introduce LayerCollapse regularization.
+The LayerCollapse regularizer works by modulating the negative slope of the ReLU non-linearity to one, effectively transforming it into a linear transformation.
+This modification not only facilitates the fusion of successive linear layers, conserving computational resources, but also can enhance model generalizability.
+This paper presents several key contributions to the field of neural network regularization and compression:
+  • We introduce LayerCollapse , a novel regularization approach that enables effective one-shot compression after training.
+  • We make a theoretical analysis providing bounds on the compression loss associated with LayerCollapse .
+  • We propose a compression-aware regularization loss designed to encourage simpler model structures during the training and finetuning phase.
+  • Through experiments across various datasets in both vision and NLP fields, we demonstrate that LayerCollapse achieves post-training compression rates of up to 74% for VGG models and 16% for transformer architectures.
+These experiments also highlight enhanced model generalization and reduced overfitting.
+  • We offer an open-source library for LayerCollapse , facilitating its easy integration with PyTorch frameworks. [^1] https://github.com/sohail2810/LayerCollapse
+2 Background
+Multi-Layer Perceptrons (MLPs) are foundational elements in modern deep learning architectures, integral to tasks such as classification, regression, and sequence modeling.
+Structurally, an MLP comprises fully connected layers where each neuron in a layer connects to every neuron in the subsequent layer.
+Originally developed as one of the first neural network models Haykin (1994), MLPs remain essential in contemporary architectures like transformers and vision-based networks Vaswani et al.
+(2017); Dosovitskiy et al.
+(2020) due to their capacity to model complex relationships and approximate intricate functions.
+2.1 Role and Structure of MLPs
+MLPs consist of sequences of linear transformations interspersed with non-linear activation functions, where the linear transformation can be represented as $\mathbf{y}=\mathbf{W}\mathbf{x}+\mathbf{b}$.
+Without the non-linear functions, consecutive linear transformations would collapse into a single transformation, limiting the model’s ability to learn non-linear relationships Goodfellow (2016).
+Popular activation functions in MLPs include ReLU, parametric ReLU (PReLU), GeLU, ELU, Sigmoid, and Tanh, with ReLU widely used for its computational efficiency and mitigation of the vanishing gradient problem Nair & Hinton (2010); Glorot et al.
+(2011).
+2.2 MLPs in Modern Architectures
+MLPs are pivotal in advanced architectures like transformers, which utilize MLP layers following self-attention mechanisms to enhance representational capacity Vaswani et al.
+(2017).
+In Vision Transformers (ViTs), MLPs process image patch embeddings, enabling spatial and feature learning from image data Dosovitskiy et al.
+(2020).
+Architectures like MLP-Mixers further showcase the versatility of MLPs by leveraging them for token and channel mixing, demonstrating that MLPs alone can achieve competitive performance in vision tasks Tolstikhin et al.
+(2021).
+2.3 Importance of Non-Linear Activations
+The inclusion of non-linear activation functions between linear layers is essential for MLPs to capture complex, non-linear mappings.
+Activation functions like ReLU introduce discontinuities that prevent the model from being purely linear, while bounded functions like Sigmoid and Tanh enable stability in certain applications despite their tendency to saturate Glorot et al.
+(2011).
+By breaking up linear layers, non-linearities enable MLPs to approximate complex functions effectively, a quality exploited in diverse applications.
+2.4 Applications and Challenges
+MLPs are critical in classification heads for converting high-dimensional representations to class scores in tasks such as image classification and sentiment analysis Russakovsky et al.
+(2015).
+In regression tasks, MLPs predict continuous values and are widely used in financial forecasting and scientific applications.
+Despite these strengths, the dense nature of MLPs contributes significantly to the computational overhead in models, spurring the need for optimization techniques like pruning, quantization, and activation modulation Fan et al.
+(2019); Hubara et al.
+(2017).
+Activation modulation, in particular, can simplify models by modulating non-linear activations toward linearity, thus reducing computational demands without heavily impacting model expressiveness.
 3 Methodology
-Figure 2: Overview of the LayerCollapse process: $(a)$ The compression aware the regularization loss shifts the PReLU towards linearity.
-$(b)$ The layer collapse operation involves eliminating a layer by replacing the hidden layer and adjusting the weights through the product of two matrices.
-This section introduces LayerCollapse, an innovative compression method that adaptively reduces the model’s depth through regularization.
-This methodology aims to optimize the size of the model based on the quality of the dataset, resulting in enhanced generalization.
-Our proposed framework seamlessly integrates regularization and compression.
-Unlike pruning [24] and quantization [15] which require the compressed model to be finetuned, LayerCollapse requires the finetuning prior to the compression phase.
-As depicted in Figure 2, during a compression aware training or finetuning, we regularize a layer’s complexity by linearizing its activation function.
-We define Collapsing an MLP on a model that has shifted at least one of its non-linear layers into linearity.
-As depicted in Figure 2 Linearizing a non-linear activation within a model allows for the collapse of the two surrounding linear layers into one linear layer, effectively reducing the depth.
-The resulting model has fewer parameters and a smaller depth, requiring less resources for execution.
-For post training compression, one needs to run a few cycles of finetuning using the LayerCollapse regularizer to ensure the linearization loss is marginal.
-We define a Collapsible MLP to be two consecutive fully connected layers utilizing a Parametric Rectified Linear Unit (PReLU) [12] in the middle.
+This section presents LayerCollapse , a novel regularization technique designed to adaptively compress the model by reducing its depth, possibly improving generalization in the process.
+In LayerCollapse , we employ compression-aware training.
+This approach enables the extraction of a shallower model directly, without the need for finetuning.
+As illustrated in Figure 2, LayerCollapse involves regularizing the complexity of a layer by linearizing its activation function when training or fine-tuning.
+This linearization enables the merging of adjacent linear layers into a single layer, thus effectively reducing the model’s depth.
+The resulting model has lower storage, and computation demands without the need for software/hardware support.
+To gain intuition on the effect that the non-linearities have on the model’s representation power and complexity, we present a toy scenario.
+Consider a 2-layer MLP with a hidden dimension of 500 and input and output dimensions of 1, employing a ReLU non-linearity.
+This network can be used to fit a polynomial function $f(x)=(x-3)(5x+2)(x+1)$ with samples having Gaussian noise.
+We apply gradient descent to fit the MLP to this function, aiming to uncover the underlying function $f(x)$.
+The results of this experiment are shown in Figure 1.
+In this demo, the regular model is described above.
+The $n\%$ linear model is a leaky ReLU with a negative slope of $n\%$.
+We observe that at $n=0.8$, the complexity of the 2-layer network is reduced such that it does not overfit to noise, closely approximating the underlying function.
+Conversely, at $n=0.98$, the reduced complexity of the model leads to underfitting.
+Setting $n=100\%$ results in a linear function, $Activation(x)=x$, effectively simplifying the 2-layer MLP to just a linear transformation since the product of the matrices $L_{1}^{1\times m}$ and $L_{2}^{m\times 1}$ results in a scalar.
+Here, $L_{1}\in\mathbb{R}^{1\times m}$ and $L_{2}\in\mathbb{R}^{m\times 1}$ are the weight matrices of the first and second layers.
+Following the intuition from this experiment, we can see the effect of the negative slope on the model’s expressiveness.
+We design a regularizer that adaptively controls the negative slope of the MLP, preventing the model from overfitting to the noise.
+If for any layer we find that the negative slope is one, we can remove that layer from the network following the procedure in Algorithm 1 without losing any performance.
+The resulting network will have a shallower depth, fewer parameters, and lower resource requirements for execution.
+3.1 Formal Definition
+We define a Collapsible MLP as: Two consecutive fully connected layers utilizing a Parametric Rectified Linear Unit (PReLU) He et al.
+(2015) as the activation.
 The Parametric Rectified Linear Unit (PReLU) can be expressed as:
@@
 $Y_{\alpha}=W_{2}(PReLU_{\alpha}(W_{1}X+b_{1}))+b_{2}$ (2)
-where, $W_{1}$ and $W_{2}$ are the weight matrices and $b_{1}$ and $b_{2}$ are the bias vectors for the first and second fully connected layers, respectively.
+where, $X$ is the input random variable and $Y$ is the output random variable.
+$W_{1}$ and $W_{2}$ are the weight matrices and $b_{1}$ and $b_{2}$ are the bias vectors for the first and second fully connected layers, respectively.
 $\alpha$ can be used to make the activation linear.
@@
 $Y_{linear}=W_{2}W_{1}X+W_{2}b_{1}+b_{2}$ (3)
-We define the error as the squared difference of $Y_{\alpha}$ and $Y_{linear}$. In Appendix A, we show that $\forall\delta>0$
+We define the non-linearity error as the squared difference of $Y_{\alpha}$ and $Y_{linear}$. In Appendix A, we show that $\forall\delta\in[0,1]$
 $P\left\{|Y_{\text{linear}}-Y_{\alpha}|^{2}\leq C\times(1-\alpha)^{2}\right\}>1-\delta$ (4)
@@
 $C=\sigma_{max}(W_{2}W_{1})^{2}|x^{\delta}|^{2}+|W_{2}b_{1}|^{2}$
-where, $\sigma_{max}(W_{2}W_{1})$ is the largest singular value of $W_{2}W_{1}$ By choosing $\delta$ arbitrarily small, error is bounded proportional to $(1-\alpha)^{2}$ with probability one.
-We capitalize on this by incorporating $(1-\alpha)^{2}$ into our loss function.
+where, $\sigma_{max}(W_{2}W_{1})$ is the largest singular value of $W_{2}W_{1}$ By choosing $\delta$ arbitrarily small, probability of error is bounded proportional to $(1-\alpha)^{2}$ with probability one.
+We incorporating $(1-\alpha)^{2}$ into our loss function to optimize for the non-linearity loss as well as the label cross-entropy loss.
 This loss term serves to align $Y_{\alpha}$ with $Y_{linear}$, compressing the model into a shallower model as demonstrated in Figure 2.
-In simpler terms, when $\alpha$ approaches one, substituting $Y_{linear}$ with $Y_{\alpha}$ with an error bounded by a constant multiplied by $(1-\alpha)^{2}$.
-In practical applications, Dropout [34] and BatchNormalization [16] are frequently used between the layers of a network.
+In simpler terms, as $\alpha$ approaches one, substituting $Y_{linear}$ with $Y_{\alpha}$ incurs a bounded error probability, the error probability is proportional to $(1-\alpha)^{2}$ times a constant.
+In practical applications, Dropout Srivastava et al.
+(2014) and BatchNormalization Ioffe & Szegedy (2015) are frequently used between the layers of a network.
 Dropout works by temporarily deactivating random neurons during training, which helps spread the learning across different neurons to avoid over-reliance on any particular subset of them.
 It is important to note that Dropout is not active during model deployment, meaning it behaves as if it is not there.
-This characteristic means that Dropout doesn’t affect the layer collapsing process and can be left out when considering it.
+This characteristic means that Dropout doesn’t affect the layer collapsing process and can be left out of consideration.
 BatchNormalization is a process described by the formula:
 $Y=\frac{X-E[X]}{\sqrt{Var[X]}}\times\gamma+\beta$ (6)
-where, $\gamma$ and $\beta$ are learnable parameters, the $E[X]$ and $Var[X]$ are the mean and variance of the current batch, which we sometimes adjust with a method called Polyak averaging to keep the batch-to-batch changes smooth.
+where, $X$ and $Y$ are the input and output random variables. $\gamma$ and $\beta$ are learnable parameters, and $E[X]$ and $Var[X]$ are the mean and variance of the current batch.
 Algorithm 1 outlines the specific steps to be taken for adjusting weights and biases when BatchNormalization is present.
-To measure how well our compression method works, we use the concept of “compression gain”.
-This is calculated as the percentage reduction in the number of parameters of a layer, excluding bias and BatchNormalization parameters.
-The formula for compression gain is:
-$Gain=1-\frac{n_{in}\times n_{out}}{n_{hidden}\times(n_{in}+n_{out})}$ (7)
-where, $n_{in}$, $n_{hidden}$, and $n_{out}$ denote the input size, hidden size, and output size of a 2-layer MLP, respectively.
-An interesting aspect of this formula is that the gain can sometimes be negative.
-Positive gain happens when $n_{hidden}$ is larger than the product of $n_{in}$ and $n_{out}$ divided by their sum.
-This means that MLPs designed with a “bottleneck” structure (where $n_{hidden}$ is relatively small) may not see much benefit from our layer collapsing method.
-On the other hand, most modern MLP designs, which tend to expand the layer size (“widening” structure), are likely to experience significant positive gains from our method.
-3.1 Adaptation for Convolutional Layers
+3.2 Compression Ratio
+Compression ratio for the collapsed layer can be calculated as:
+$CR=1-\frac{n_{in}\times n_{out}}{n_{hidden}\times(n_{in}+n_{out})}$ (7)
+where $CR$ is the compression ratio, $n_{in}$, $n_{hidden}$, and $n_{out}$ denote the input size, hidden size, and output size of the 2-layer MLP, respectively.
+Noteworthy is that the compression ratio can be negative.
+Positive compression ratio only happens when $n_{hidden}$ is larger than the product of $n_{in}$ and $n_{out}$ divided by their sum.
+This means that MLPs designed with a “bottleneck” structure (where $n_{hidden}$ is relatively small) may not see a benefit from LayerCollapse .
+However, most modern MLP designs tend to expand the layer size (“widening” structure), and are likely to have significant compression ratios.
 Convolutional layers play a pivotal role in numerous neural network architectures.
-Essentially convolution can be formulated as sparse fully connected layers with some redundancies.
+Essentially, convolution can be formulated as sparse fully connected layers with some redundancies.
 Their nature as linear transforms permits representation via circulant matrices.
 A circulant matrix is a matrix where the row vectors have the same elements and each row vector is rotated one element to the right relative to the preceding row vector.
-An interesting property of these matrices is that their product also forms a circulant matrix, resulting in another convolution layer.
-The kernel size of the resulting convolution layer, denoted as $k$, is calculated by $k=k_{1}+k_{2}-1$ where, $k_{1}$ and $k_{2}$ are the kernel sizes of the original convolution layers.
-The compression gain can be calculated by:
-$Gain_{conv}=1-\frac{k_{1}^{2}\times c_{in}\times c_{hidden}+k_{2}^{2}\times c_{hidden}\times c_{out}}{(k_{1}+k_{2}-1)^{2}\times c_{in}\times c_{out}}$ (8)
+The product of two circulant matrices is also a circulant matrix.
+So collapsing two convolution layers will result in another convolution layer.
+The kernel size of the resulting convolution layer, denoted as $k$, is calculated as $k=k_{1}+k_{2}-1$ where, $k_{1}$ and $k_{2}$ are the kernel sizes of the original convolution layers.
+Compression gain for convolution layers can be calculated as:
+$CR_{conv}=1-\frac{k_{1}^{2}\times c_{in}\times c_{hidden}+k_{2}^{2}\times c_{hidden}\times c_{out}}{(k_{1}+k_{2}-1)^{2}\times c_{in}\times c_{out}}$ (8)
 where, $c_{in}$, $c_{hidden}$, and $k_{1}$ correspond to the input channels, output channels, and the kernel size of the initial convolution layer, respectively.
-Meanwhile, $c_{hidden}$, $c_{out}$ and $k_{2}$ are the parameters of the subsequent convolution layer.
-However, an essential note when applying LayerCollapse to common CNN architectures, such as ResNet, is that the gain often turns out to be negative.
-To achieve a positive gain, it is imperative for $c_{hidden}$ to be significantly larger than both $c_{in}$ and $c_{out}$.
-This observation highlights the need for careful configuration when attempting to collapse layers in such architectures.
-Overall, this framework offers a principled approach to explore the potential benefits of LayerCollapse on convolution layers, illuminating the optimization potential for convolutional neural networks.
-3.2 Prior Distribution Regularization
-We extend the use of the parametric ReLU (PReLU) function to develop a regularization method based on the maximum a posteriori (MAP) framework.
-Aligning with Occam’s razor, which favors simpler explanations, we propose to adopt a normal distribution with a mean of 1 as the prior distribution for $\alpha$.
-This setting effectively limits the two layers to a linear space, reducing their functional complexity, as illustrated in Figure 1.
+$c_{hidden}$, $c_{out}$ and $k_{2}$ are the parameters of the subsequent convolution layer.
+However, a crucial point to consider when applying LayerCollapse to common CNN architectures, such as ResNet, is that the compression ratio is negative.
+This is due to the fact that a positive compression ratio requires for $c_{hidden}$ to be significantly larger than both $c_{in}$ and $c_{out}$.
+This observation highlights the need for careful consideration when attempting to collapse layers in such architectures.
+3.3 Prior Distribution Analysis of Regularization
+We extend the use of the parametric ReLU (PReLU) function to develop a regularization method based on the maximum a posteriori (MAP) framework DeGroot (2005).
+Aligning with Occam’s razor Schaffer (2015), which favors simpler explanations, we propose to adopt a normal distribution with a mean of 1 as the prior distribution for $\alpha$.
+This setting effectively limits the two layers to a linear space, reducing their functional complexity.
 This choice is backed by empirical results given in section 4.
-MAP framework suggests adding negative log likelihood of the prior distribution of parameters to the loss function.
+An intuition for this choice is presented in figure 1 MAP framework suggests adding negative log likelihood of the prior distribution of parameters to the loss function.
 This is referred to as the prior loss or regularization loss.
@@
 $\mathcal{L}_{\text{reg}}=lc\times(1-\alpha)^{2},\text{}$ (9)
-where $lc$ is the regularization constant.
+where $lc$ is the regularization constant. Notice this formulation aligns perfectly to our intuition from equation (4).
 Algorithm 1 Layer Collapse
-1:function Collapse($W_{1},b_{1},W_{2},b_{2},\gamma,\beta,\bar{\mu},\bar{\sigma},\alpha,\tau$)
-2:     if $\alpha\leq\tau$ then
-3:         $s_{BN}\leftarrow\frac{\gamma}{\sqrt{\bar{\sigma}}}$
-4:         $W_{\text{new}}\leftarrow s_{BN}\times W_{2}\cdot W_{1}$
-5:         $b_{\text{new}}\leftarrow W_{2}\cdot(s_{BN}\times(b_{1}-\bar{\mu})+\beta)+b_{2}$
-6:         return $W_{\text{new}},b_{\text{new}}$
-7:     else
-8:         return un-collapsible
-9:     end if
-10:end function
+  function Collapse($W_{1},b_{1},W_{2},b_{2},\gamma,\beta,\bar{\mu},\bar{\sigma},\alpha,\tau$
+     if $\alpha\leq\tau$ then
+        $s_{BN}\leftarrow\frac{\gamma}{\sqrt{\bar{\sigma}}}$
+        $W_{\text{new}}\leftarrow s_{BN}\times W_{2}\cdot W_{1}$
+        $b_{\text{new}}\leftarrow W_{2}\cdot(s_{BN}\times(b_{1}-\bar{\mu})+\beta)+b_{2}$
+        return $W_{\text{new}},b_{\text{new}}$
+     else
+        return un-collapsible
+     end if
+  end function
+Table 2: Training performance on Wikitext-103 dataset. For base variants the modified layer is the last layer i.e. layer 12. In the large variant layers 11, 16, and 32 are selected for compression.
+Model Layers Params (M) PPL
+GPT-2 Radford et al. (2019) 0 124.4 18.87
+GPT-2 + LayerDrop Fan et al. (2019) 1 119.9(-4%) 19.29
+GPT-2 + LayerCollapse (ours) 1 120.3(-3%) 19.09
+GPT-2 Large Radford et al. (2019) 0 774.0 17.65
+GPT-2 Large + LayerDrop Fan et al. (2019) 3 734.7(-5%) 15.61
+GPT-2 Large + LayerCollapse (ours) 3 739.6(-4%) 15.17
+3.4 Alternative Activation Functions
+Thus far, we have focused on the ReLU function as a primary non-linear activation function.
+ReLU’s simplicity and efficiency have had a significant impact on deep learning since its introduction.
+Recently, several ReLU variants, including ELU Clevert (2015), GeLU Hendrycks & Gimpel (2016), and SiLU Elfwing et al.
+(2018), have gained popularity due to their enhanced flexibility.
+We propose a general approach to smoothly linearize ReLU-family activations.
+For functions like ELU, which are linear in the positive region and non-linear in the negative, we incorporate a linear term, $y=x$, into the negative region.
+This term is blended with the original function using a weighted sum, where the weight parameter is denoted by $\alpha$ as defined in section 3.1.
+For example, the modified ELU function can be written as:
+$\text{ELU}^{\beta}(x)=\begin{cases}x&\text{if }x>0\\ \alpha x+(1-\alpha)\beta(\exp(x)-1)&\text{if }x\leq 0\end{cases}$
+where $\beta$ is the ELU parameter. When $\alpha=1$, ELU transitions to a linear function. To ensure $\alpha$ remains within $[0,1]$ during training, we apply a sigmoid to it.
+For one-region activations like GeLU and SiLU, which follow the form $f(x)=x\times h(x)$ with $h(x)\rightarrow 1$ as $x\rightarrow\infty$ and $h(x)\rightarrow 0$ as $x\rightarrow-\infty$, we adjust using $\alpha$ in the form $f(x)=x\times(h(x)+\alpha(1-h(x)))$, where $\alpha\in[0,1]$.
+This leads to parametric versions:
+$\text{GeLU}^{\alpha}(x)=x(\Phi(x)+\alpha(1-\Phi(x)))$
+where $\Phi(x)$ is the Gaussian cumulative distribution function, and for SiLU:
+$\text{SiLU}^{\alpha}(x)=x(\sigma(x)+\alpha(1-\sigma(x)))$
+where $\sigma(x)=\frac{e^{x}}{1+e^{x}}$ is the sigmoid function.
 4 Experiments
-In this section we evaluate our method on both post training compression, and compression aware training with our regularization.
-Initially, we explore the standalone capability of LayerCollapse in post training compression of streamlined vision models including vision transformer, mlp-mixer and VGG.
-We then compare LayerCollapse with Knowledge Distillation on the cost of training and achievable results on a target network.
-Subsequently, we demonstrate the effectiveness of compression aware training in an ablation study on mid-sized datasets: TinyImageNet [22], and CIFAR100 [21].
-Further extending the scope of our investigation, we exhibit preliminary results on the BERT [2] language model to provide a proof-of-concept for the generalizability of LayerCollapse in domains beyond computer vision, such as natural language processing (NLP).
-These experiments aim to demonstrate LayerCollapse’s versatility and its potential to enhance efficiency and performance in various machine learning contexts.
-Model Collapsed Layers No. of Params (M) Top1 Acc. (%) Top5 Acc. (%) MACs (G)
-ViT T/16 0 5.72 75.35 93 1.08
-1 5.46 75.79 93 1.02
-2 5.20 75.35 92 0.97
-3 4.94 74.42 92 0.91
-ViT S/16 0 22.05 81.38 96 4.24
-1 21.02 80.96 96 4.04
-2 19.98 79.74 95 3.84
-ViT B/16 0 86.57 81.43 96 16.85
-1 82.44 82.17 96 16.04
-2 78.3 80.31 95 15.23
-ViT L/16 0 304.33 84.80 98 59.66
-2 289.64 84.22 97 56.77
-Mixer B/16 0 59.88 76.47 92 12.61
-2 55.64 76.11 92 11.71
-4 51.39 75.67 91 10.81
-VGG19 0 143.67 74.21 92 19.65
-2 45.11 71.22 88 19.55
-VGG16 0 138.36 73.36 92 15.48
-2 39.80 67.16 88 15.39
-VGG13 0 133.05 71.58 90 11.32
-2 34.49 67.21 88 11.22
-VGG11 0 132.86 70.37 90 7.62
-2 34.31 66.36 88 7.52
-Table 2: Post training compression results, showcasing accuracy and compression attained via LayerCollapse across different layers in relevant architectures.
-4.1 Compression
-Our compression experiments are performed on ImageNet-1k classification task [30].
-Unlike pruning [24] and quantization [15] compression methods which require finetuning on the compressed method to recover accuracy, LayerCollapse utilizes finetuning with the regularizer to ensure accuracy.
-Afterwards the compression phase will endure no accuracy loss.
-We first assess the stand-alone compression efficacy of LayerCollapse.
-This is followed by a comparative analysis of LayerCollapse and Knowledge Distillation, focusing on training efficiency and resource utilization.
-4.1.1 Stand-Alone Compression
-Dataset. We employ the ImageNet-1k [30] dataset for both fine-tuning and validation processes. It comprises over 1.2 million images for training and 50,000 for validation.
-Implementation.
-We adapt pre-trained models from the TIMM library [39] for MLP-Mixer [35] and ViT [7], alongside VGG variants from torchvision [28].
-ViT and Mixer architectures use GeLU activations in their original implementation.
-In target layers, these are converted to ReLU to align with the requirements of LayerCollapse.
-The fine-tuning follows a consistent procedure across models, using Stochastic Gradient Descent (SGD) with momentum and a learning rate of $5e^{-4}$ as the optimizer.
-The loss function is comprised of three parts, cross entropy between labels and result, KL divergence between the output and uncollapsed models output and the LayerCollapse regularization term (9) with $lc=0.2$.
-This process is capped at 20 epochs.
-Each layer for all ViT experiments use 5 epochs.
-Number of epochs for VGG architectures are given in Table 3.
-The MAC operations are all calculated by propagating a random $224\times 224$ image through the model.
-Further details are enumerated in Appendix B.
-Sensitivity Analysis.
-To quantify the robustness of LayerCollapse, we conduct a layer-by-layer sensitivity analysis on the Vision Transformer T/16 model, as shown in Figure 3.
-This removal of layers provides insight into the impact of incremental compression on the model’s accuracy.
-Interestingly, we observe an initial increase in accuracy upon the collapse of the first layer, indicating an improvement in generalization compared to the original model.
-Figure 3: Layer-wise collapse accuracy and parameter reduction analysis for ViT-T/16.
-Model Method Top1 Acc(%) Top5 Acc(%) No. of Parameters (M) Cost Epochs
-Initial Final
-vgg19 LC 71.22 88 143.67 45.11 11h 5
-KD 63.42 85 51h 23
-vgg16 LC 70.01 89 138.36 39.80 14h 7
-KD 59.63 82 31h 16
-vgg13 LC 67.21 88 133.05 34.49 16h 10
-KD 58.26 81 45h 28
-vgg11 LC 66.36 88 132.86 34.31 20h 20
-KD 58.64 81 42h 42
-Table 3: A comparative analysis of Layer Collapse (LC) and Knowledge Distillation (KD) methods applied to VGG models.
-Discussion.
-Table 2 illustrates that LayerCollapse achieves a substantial reduction in model size—up to 74% in VGG11—and computational demands with minimal accuracy degradation.
-Notably, models with heavier classification heads, such as VGG, can benefit significantly from our method.
-ViT models, the T/16 and B/16 variants, get an increase in accuracy with the first layer collapse.
-This can be attributed to an increase of generalization in simpler models.
-In ViT T/16, S/16, and B/16 we get roughly 5 percent reduction in model size per LayerCollapse.
-In all ViT variants we get a consistent reduction of over 10% in size and complexity with less than a 1% accuracy loss.
-In the Mixer architecture, 2 types of MLP layers known as MLP-token and MLP Channels exist.
-MLP-Channel has a significantly bigger size.
-We found that MLP-tokens can be collapsed in about an epoch of computation whilst MLP-channels require more effort.
-In the Mixer experiment, we collapsed the layers in pairs to ensure MLP-Channels are affected.
-In the Mixer architecture we get 14% reduction in size with less than a 1% accuracy degradation.
-Notice that the Mixer architecture is the only one that gets a linear reduction in MAC operations with each pair collapsed.
-We used the idea first introduced in Knowledge Distillation to enhance the performance of our finetuning procedure.
-Using the KL-Divergence loss, we speedup the gradient flow and increase the convergence speed.
-4.2 LayerCollapse vs Knowledge Distillation
+We conduct a comprehensive evaluation of LayerCollapse ’s accuracy and efficiency.
+For accuracy, we demonstrate LayerCollapse ’s capability on NLP and vision tasks, showcasing its effectiveness as both a regularizer and a compression method.
+Experiments presented in Section 4.1 illustrate LayerCollapse ’s dual role not only as a compression tool for pre-trained networks but also in enhancing model generalization.
+We compare LayerCollapse against a modified version of LayerDrop, a structured pruning technique, highlighting the distinct advantages of our approach.
+Detailed in Section 4.2, we finetune models with LayerCollapse for regularization followed by compression.
+Additionally, we compare LayerCollapse against Knowledge Distillation, as discussed in Hinton et al.
+(2015), focusing on the computational resources required for each method, despite targeting the same network architectures.
+Finally, we test LayerCollapse as a regularizer by training a VGG network from scratch, presenting two scenarios: using LayerCollapse and not using it, to evaluate its impact on model performance and efficiency.
+Table 3: Post-training compression GLUE test results illustrating accuracy and achieved compression.
+For QQP, MRPC average of F1 and accuracy are reported, for STS-B average of Spearman and Pearson is reported.
+Model Layers Params (M) MNLI(m/mm) SST-2 STS-B RTE QNLI QQP MRPC CoLA Average
+Bert Base 0 109.1 83.5/83.7 92.7 86.4 65.7 91.0 89.4 85.8 61.0 81.9
+4 92.5(-15%) 82.6/82.8 91.2 87.3 61.4 89.8 86.3 88.4 59.6 80.8
+Bert Large 0 334.9 86.3/86.1 93.5 87.7 66.4 89.3 89.9 85.8 62.6 82.7
+4 309.7(-8%) 86.6/85.5 93.0 89.5 71.5 91.4 89.4 87.5 62.6 83.8
+4.1 Regularization Performance of LayerCollapse
+4.1.1 Pre-Trained Models
 Setup.
-Our evaluation framework utilizes batch-normalized VGG variants from the torchvision repository [28], with ImageNet-1K [30] providing a substantial dataset for training and validation.
-These variants are chosen for their prevalent use in compression studies and their capability to learn without prior training.
-To facilitate a fair comparison between LayerCollapse and Knowledge Distillation, we adopt the same target architecture for both methods.
-The rest of the training is set up identical to section 4.1.1 for both LayerCollapse and Knowledge distillation.
-Comparison.
-The models are rigorously compared across several dimensions: training cost, the number of training epochs required, and the accuracy metrics—top-1 and top-5—on the validation set.
-The computational costs are quantified on an Nvidia RTX A6000 GPU.
-Discussion.
-Referencing Table 3, LayerCollapse consistently achieves over 10% higher accuracy while incurring only 30% of the computational cost compared to Knowledge Distillation.
-This trend is more pronounced in larger models, which exhibit greater compression potential, and result in up to 15% accuracy gain with 50% of tuning cost.
-It is noteworthy that for models necessitating pre-training, such as MLP-Mixer and ViT, the absence of pre-training data and the high computational demands often make Knowledge Distillation impractical.
-LayerCollapse emerges as a unique method for model compression through architectural modification, offering significant cost advantages as demonstrated in Table 3.
-However, it is important to recognize the inherent limitations of LayerCollapse in terms of the diversity of achievable target model architectures.
-4.3 Ablation Study of Regularization
+In our experiments, we utilize the base and large versions of the GPT-2 causal language model Radford et al.
+(2019), with a fixed context size of 128 tokens.
+Our layer selection methodology, inspired by Fan et al.
+(2019), involves initially collapsing each layer, using the algorithm presented in 1, in a one-shot manner, followed by an evaluation to identify layers most tolerant to compression.
+We evaluate LayerCollapse on the widely used perplexity metric on the test set of the Wikitext-103 dataset Merity et al.
+(2016).
+We conduct training for $3k$ steps, maintaining a batch size of 128 across all experiments.
+The learning rate scheduling is managed by LiveTune Shabgahi et al.
+(2023), starting from an initial rate of $0.001$.
+Optimization is performed using a modified version of the Adam optimizer Kingma & Ba (2014), with comprehensive details about this optimizer available in Appendix B.
+An exhaustive list of experiment details can be found in Appendix C.
+The comparison was done with a modified version of LayerDrop Fan et al.
+(2019).
+This modification, drawing inspiration from Savarese & Figueiredo (2017), employs gates for selected layers.
+These gates use a sigmoid function to dynamically alternate between activation and deactivation, enhancing training stability.
+Results.
+As shown in Table 2, LayerCollapse surpasses LayerDrop in terms of final perplexity score.
+Notably, the large variant of GPT-2 tends to overfit, resulting in higher Perplexity compared to both regularized approaches.
+However, LayerDrop achieves greater compression gain.
+4.1.2 Trained from Scratch.
 We provide an in-depth analysis of how our regularization approach impacts both the generalization and compression abilities in training image classifiers from scratch.
 Dataset.
-Our study utilizes the CIFAR100 [10] and TinyImageNet [22] datasets, which are mid-sized and rich in diversity.
+Our study utilizes the CIFAR100 Han et al.
+(2015) and TinyImageNet Le & Yang (2015) datasets, which are mid-sized and rich in diversity.
 The CIFAR100 dataset includes 60,000 images across 100 classes, and TinyImageNet, a subset of the larger ImageNet, comprises 100,000 images across 200 classes.
@@
 Note that the goal of this study is not to show state of the art performance using our proposed regularization, but to show its effectiveness in reducing over-fitting and final model size.
-In all experiments, traditional regularizers, such as data augmentation (rotation, random crop, and flipping), dropout[5], and batch normalization[16], have been used.
+In all experiments, traditional regularizers, such as data augmentation (rotation, random crop, and flipping), dropoutdo Santos et al.
+(2021), and batch normalizationIoffe & Szegedy (2015), have been used.
 This is to show the added contribution of 9.
@@
 The regularization is applied to 40% of the layers, starting with the latest layers, with the strength of 0.05 across all experiments.
-extensive experiment details can be found in Appendix B.
+extensive experiment details can be found in Appendix C.4.
 Discussion.
@@
 The learning curve presented in Figure 3 suggests a narrowing gap between training and validation accuracy, indicating a reduction in overfitting and highlighting the effectiveness of our regularization strategy in improving generalization.
-4.4 LayerCollapse in NLP
-We extended LayerCollapse to the NLP domain. This section presents our initial findings using the BERT-Base model [2], a well-established pre-trained NLP model.
-Setup In our evaluation we use the General Language Understanding Evaluation on the Microsoft Research Paraphrase Corpus (GLUE-MRPC)[37][6].
-This benchmark assesses a model’s capacity to discern paraphrased sentences, determining if they are equivalent.
-Further details are found in Appendix B
-Discussion We report our results in Table 5.
-These include accuracy metrics and the number of parameters for each layer of BERT-Base modified by LayerCollapse.
-The experiment shows a promising 19% reduction in parameters with a minimal accuracy loss of less than 3%.
-Notably, we exclusively use the data in the GLUE-MRPC task in the LayerCollapse procedure.
-These preliminary results in the NLP sector suggest the versatility and potential of LayerCollapse across different domains, not limited to computer vision.
-They point to the method’s adaptability and effectiveness in optimizing model parameters while preserving accuracy without the need for pre-train dataset.
 Figure 3: Evaluating the impact of LayerCollapse regularization on CIFAR100, we compare the top-1 accuracy of VGG11 both with and without this regularization technique.
@@
 VGG19(LC) $20.1$M $63.8$ $20.4$M $48.2$
-Table 5: Layer-wise analysis of accuracy and parameters for Bert-Base via LC
-Layers Accuracy (%) Params (M)
-0 $85.0$ $109.40$
-1 $85.8$ $105.35$
-2 $84.1$ $101.20$
-3 $84.3$ $97.10$
-4 $82.1$ $92.95$
-5 $82.3$ $88.80$
-6 $82.3$ $84.70$
-7 $80.4$ $80.56$
-8 $79.4$ $76.40$
-5 Conclusion
-This paper introduces LayerCollapse, a novel method for compressing neural networks through regularization.
-LayerCollapse simplifies a network by merging pairs of linear layers, reducing the network’s depth, the number of parameters, thereby, reducing operational requirements.
-We introduce a compression aware regularizer that encourages the development of shallower network architectures.
-Under rigorous evaluation, LayerCollapse demonstrates state-of-the-art performance in compressing Vision Transformers, improving even improving accuracy in some cases.
-In comparison to Knowledge Distillation, LayerCollapse consistently demonstrated superior efficiency in both computation and final results, achieving over 10% better performance while utilizing 80% fewer resources.
-LayerCollapse marks a significant paradigm shift in efficient neural network design and bridges the gap between performance and model size.
-Supplementary Material
+4.2 LayerCollapse for Compression
+We apply LayerCollapse to compress targeted layers in pretrained models in both computer vision and NLP tasks.
+To utilize LayerCollapse as compression for pre-trained models, a few rounds of fine-tuning with the LayerCollapse regularizer are required to ensure the linearization does not significantly impact accuracy.
+The process involves a few training steps with the regularization loss defined in Equation 9.
+In all cases, we set the slope threshold for layer collapse at $0.0001$.
+It’s important to note that we report results directly after layer collapse, without the need for additional fine-tuning.
+For architectures such as Bert, ViT, and MLP-Mixer, we substitute GeLU activations in the compressed layers with ReLU, as GeLU is a smoothed ReLU and this change has been found to be insignificant in our tests.
+Our compression approach focuses on the model’s final layers, driven by the insight that these layers handle more abstract data representations, rendering them more suitable for compression.
+The compression is carried out one layer at a time.
+For detailed implementations notes, please refer to Appendix C.
+4.2.1 Performance on GLUE
+Setup.
+We assess LayerCollapse ’s compression effectiveness on pre-trained Bert-Large and Bert-Base Devlin et al.
+(2018) using the GLUE benchmark Wang et al.
+(2018).The Average column is slightly different than GLUE score since we exclude WNLI as discussed in Devlin et al.
+(2018).
+Results.
+As illustrated in table 3 LayerCollapse yields $15\%$ compression gain for Bert base with about $1\%$ score reduction on the GLUE benchmark.
+As evident the larger variant of Bert is prone to overfit thus LayerCollapse regularization improves the score in many tasks.
+As illustrated LayerCollapse provided better scores than the baseline in Devlin et al.
+(2018).
+This is while the resulting model is $8\%$ smaller in terms of parameter count.
+Table 5: Post-training compression results, showcasing accuracy and compression attained via LayerCollapse across different layers in relevant architectures.
+Model Layers Params (M) MACs (G) Top1 Acc. (%)
+ViT T/16 0 5.72 1.08 75.35
+3 4.94(-14%) 0.91(-16%) 74.42
+ViT S/16 0 22.05 4.24 81.38
+2 19.98(-9%) 3.84(-9%) 79.74
+ViT B/16 0 86.57 16.85 81.43
+2 78.30(-10%) 15.23(-10%) 80.31
+ViT L/16 0 304.33 59.66 84.80
+2 289.64(-5%) 56.77(-5%) 84.22
+Mixer B/16 0 59.88 12.61 76.47
+4 51.39(-14%) 10.81(-14%) 75.67
+VGG19 0 143.67 19.65 74.21
+2 45.11(-67%) 19.55(-1%) 71.22
+VGG11 0 132.86 7.62 70.37
+2 34.31(-74%) 7.52(-1%) 66.36
+4.2.2 Performance on ImageNet-1K
+Setup.
+We apply LayerCollapse across a diverse set of vision models, including CNNs, MLP-Mixers Tolstikhin et al.
+(2021), and Vision Transformers Dosovitskiy et al.
+(2020), utilizing the ImageNet-1K benchmark Russakovsky et al.
+(2015) for evaluation.
+Pretrained models are sourced from the timm library Wightman (2019).
+Optimization is performed using SGD with momentum, with a learning rate set to $5e^{-4}$.
+We employ the Self Distillation method Fan et al.
+(2019).
+The loss function is comprised of three parts, cross entropy between labels and result, KL divergence between the output and original model output as the teacher and the LayerCollapse regularization term given in equation (9) with $lc=0.2$.
+This process is capped at 20 epochs.
+Details are enumerated in Appendix C.
+Sensitivity Analysis.
+To quantify the robustness of LayerCollapse, we conduct a layer-by-layer sensitivity analysis on the Vision Transformer T/16 model, as shown in Figure 4.
+This removal of layers provides insight into the impact of incremental compression on the model’s accuracy.
+We observe an initial increase in accuracy upon the collapse of the first layer, attributed to an improvement in generalization compared to the original model.
+Figure 4: Layer-wise collapse accuracy and parameter reduction analysis for ViT-T/16.
+Results.
+Table 5 illustrates that LayerCollapse achieves a substantial reduction in model size-up to $74\%$ in VGG11—with minimal accuracy degradation.
+Notably, models with heavier classification heads, such as VGG, can benefit significantly from our method.
+In all ViT variants we get a consistent reduction of over $10\%$ in size and complexity with less than a 1% accuracy loss.
+In the Mixer architecture, 2 types of MLP layers known as MLP-token and MLP Channels exist.
+MLP-Channel has a significantly bigger size.
+We found that MLP-tokens can be collapsed in about an epoch of computation whilst MLP-channels require more effort.
+In the Mixer experiment, we collapsed the layers in pairs to ensure MLP-Channels are affected.
+In the Mixer architecture we get $14\%$ reduction in size with less than a $1\%$ accuracy degradation.
+4.3 LayerCollapse vs Knowledge Distillation
+Setup.
+Our evaluation framework utilizes batch-normalized VGG variants from the torchvision repository maintainers & contributors (2016).
+We use the ImageNet-1K Russakovsky et al.
+(2015) dataset, providing a substantial dataset for training and validation.
+These variants are chosen for their prevalent use in compression studies.
+To facilitate a fair comparison between LayerCollapse and Knowledge Distillation, we adopt the same target architecture for both methods.
+The rest of the training is set up identical to section 4.2 for both LayerCollapse and Knowledge distillation.
+Comparison.
+The models are rigorously compared across several dimensions: training cost, the number of training epochs required, and the accuracy metrics—top-1 and top-5—on the validation set.
+The computational costs are quantified on one Nvidia RTX A6000 GPU.
+Results.
+Referencing table 6, LayerCollapse consistently achieves over 10% higher accuracy while incurring only 30% of the computational cost compared to Knowledge Distillation.
+This trend is more pronounced in larger variants, which exhibit greater compression potential.
+Showing up to 15% accuracy gain with 50% of tuning cost.
+It is noteworthy that for models necessitating pre-training, such as MLP-Mixer and ViT, the absence of pre-training data and the high computational demands often make Knowledge Distillation impractical.
+However, we recognize the inherent limitations of LayerCollapse in terms of the diversity of achievable target model architectures.
+LayerCollapse can only remove fully connected layers while model distillation can change the architecture completely.
+5 Discussion
+In this study, we presented LayerCollapse , a novel approach for compressing neural networks through regularization techniques that strategically merge linear layer pairs, effectively reducing both network depth and parameter count.
+This approach specifically encourages the formation of shallower network structures, contributing to computational efficiency.
+Evaluations show that LayerCollapse achieves better performance in regularization and compression than other structural modifications.
+Compared to Knowledge Distillation, LayerCollapse delivers over 10% higher performance while using 80% less computational resources.
+However, the method’s strengths primarily benefit architectures featuring MLPs with widening designs, and it does not substantially improve compression for current CNN models.
+Table 6: An analysis of LayerCollapse (LC) and Knowledge Distillation (KD) methods applied to VGG variants.
+Model Method Top1 Acc(%) No. of Parameters (M) Cost Epochs
+vgg19 LC 71.22 45.11 11h 5
+KD 63.42 45.11 51h 23
+vgg16 LC 70.01 39.80 14h 7
+KD 59.63 39.80 31h 16
+vgg13 LC 67.21 34.49 16h 10
+KD 58.26 34.49 45h 28
+vgg11 LC 66.36 34.31 20h 20
+KD 58.64 34.31 42h 42
+6 Related Work
+In overparameterized neural networks, two major concerns are generalizability and computational requirements.
+Compression techniques aim to address the second issue by reducing computational demands while maintaining model precision.
+The predominant methods for model compression include pruning Han et al.
+(2015), quantization Hubara et al.
+(2017), and knowledge distillation Hinton et al.
+(2015), each contributing uniquely to model optimization Gupta & Agrawal (2022).
+Quantization reduces the number of bits needed to store weights, though it requires hardware support to be effective.
+Pruning strategies fall into two categories, structured and unstructured.
+Unstructured pruning reduces storage needs by utilizing sparse representations, but high prune rates and hardware support are necessary.
+Structured pruning Fan et al.
+(2019) Lemaire et al.
+(2019) Fu et al.
+(2022) changes the model architecture, obviating the need for additional hardware or software support.
+Knowledge distillation changes the model’s architecture by transferring knowledge from a large model to a smaller one.
+The problem with knowledge distillation is the immense computational demands of training the target architecture.
+LayerCollapse , similar to structured pruning and knowledge distillation, modifies the neural network architecture to achieve compression without requiring further software or hardware optimization.
+Our method compares favorably to LayerDrop Fan et al.
+(2019) and knowledge distillation Hinton et al.
+(2015), as shown in Section 4.
+Pruning and quantization can be iterative or one-shot.
+Iterative methods compress the model in steps with finetuning, while one-shot methods incorporate regularization during training to enable compression without further finetuning.
+Quantization-aware training (QAT) Nguyen et al.
+(2020) and pruning-aware training (PAT) Lemaire et al.
+(2019)Jiang et al.
+(2023)Wen et al.
+(2016) Alvarez & Salzmann (2016) are common practices to mitigate accuracy loss.
+LayerDrop is a notable PAT that improves training results as a regularizer Fan et al.
+(2019).
+In Section 4.1, we evaluate the regularization effect of LayerCollapse against LayerDrop.
+Explicit regularization methods, such as Dropout do Santos et al.
+(2021), Ghiasi et al.
+(2018), Srivastava et al.
+(2014), and weight decay, follow the principle that simpler models are more robust Schaffer (2015) and less prone to overfitting.
+Following the same principle, we adjust the model’s structure with LayerCollapse to prevent overfitting.
+However, this principle has been shown to be inaccurate in recent over-parameterized models when employed with immense amounts of data.
+Previous work on private inference has explored eliminating non-linearities in neural networks for computational efficiency of private inference.
+DeepReDuce Jha et al.
+(2021) and Selective Network Linearization (SNL) Cho et al.
+(2022) adopt regularization strategies for linearization.
+Unlike SNL, LayerCollapse introduces a theoretical regularizer and demonstrates significant efficiency gains across computer vision and natural language processing domains.
+7 Conclusion and Limitations
+LayerCollapse advances the development of efficient neural networks by optimizing performance alongside a reduction in model size and computational needs.
+However, the utility of LayerCollapse is somewhat restricted as it primarily compresses fully connected layers, which limits its effectiveness in CNN architectures.
+Moreover, this method necessitates either finetuning or pretraining to prepare the model for compression, and it also demands precise adjustment of the hyperparameter $\alpha$ to prevent a decline in performance and accuracy.
+These limitations highlight areas for further research to broaden the application of LayerCollapse to a more diverse array of network structures and model types.
 Appendix A Proof of Equation (4)
@@
 $P\left\{|Y_{\text{linear}}-Y_{\alpha}|^{2}\leq C\times(1-\alpha)^{2}\right\}>1-\delta$ (20)
-Appendix B Experiment Details
+Appendix B AdamLC
+Inspired by AdamW Loshchilov & Hutter (2017) we decouple the adam optimizer with our regularization so the adaptive learning rate does not make the regularization un-predictable the algorithm is given in 2.
+In this algorithm $lc$ is the regularization constant.
+Algorithm 2 AdamLC Optimization Algorithm
+0:  $\alpha,\beta_{1},\beta_{2},\epsilon,\theta_{0},lc$
+0:  Optimized parameters $\theta$
+1:  Initialize time step $t\leftarrow 0$
+2:  Initialize first moment vector $m_{0}\leftarrow\mathbf{0}$
+3:  Initialize second moment vector $v_{0}\leftarrow\mathbf{0}$
+4:  Initialize parameter vector $\theta\leftarrow\theta_{0}$
+5:  while not converged do
+6:     $t\leftarrow t+1$
+7:     $g_{t}\leftarrow\text{gradient at }\theta$
+8:     $m_{t}\leftarrow\beta_{1}\cdot m_{t-1}+(1-\beta_{1})\cdot g_{t}$
+9:     $v_{t}\leftarrow\beta_{2}\cdot v_{t-1}+(1-\beta_{2})\cdot g_{t}^{2}$
+10:     $\hat{m}_{t}\leftarrow\frac{m_{t}}{1-\beta_{1}^{t}}$
+11:     $\hat{v}_{t}\leftarrow\frac{v_{t}}{1-\beta_{2}^{t}}$
+12:     $\theta\leftarrow\theta-\alpha\cdot\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$
+13:     if $\theta$ has a negative slope then
+14:        $\theta\leftarrow\theta-2\cdot(\theta-1)\cdot lc\cdot\alpha$
+15:     end if
+16:  end while
+17:  return $\theta$
+Appendix C Experiment Details
 The experimental files are accessible at [https://github.com/sohail2810/LayerCollapse].
-B.1 Stand-Alone Compression
+C.1 Regularization Performance of LayerCollapse
+Training was conducted using a modified version of the Adam optimizer, as detailed in Appendix B.
+A consistent regularization parameter, Regularization 5, was applied across all experiments, and the same value was used for comparisons with the LayerDrop method.
+The learning rate scheduler LiveTune Shabgahi et al.
+(2023) was employed, with an initial learning rate set to $0.001$.
+In all experiments, a batch size of 128 was maintained, and the models were trained for a total of 3,000 steps.
+For performance evaluation, we applied a stride of 32 on the complete test dataset of Wikitext-103.
+C.2 LayerCollapse for Compression
+C.2.1 Performance on GLUE
+Because of the diversity of tasks each have their own requirements for training (i.e.
+optimizer, learning rate, and epochs) All tasks except for QQP the SGD optimizer is used, and the modified adamLC B is used for QQP.
+In all experiments initial learning rate is 0.0001 which reduces to 0.00001 linearly.
+Collapsing is done one layer at a time.
+Starting from the final un-collapsed layer.
+In each step we reduce the regularization loss of the active layer during a retraining phase, when the loss is bellow $1e^{-4}$ we move to the next layer.
+All tasks for Bert-base use batch size 128.
+Bert-large is trained on batch size 20.
+C.2.2 Performance of ImageNet-1k
 Dataset: We utilized the full ImageNet-1k dataset, with image dimensions of $224\times 224$.
@@
 The batch size was fixed at 256 for all experiments.
-Training Setup: We accelerated finetuning by using the original model’s predicted labels, akin to Knowledge Distillation techniques. The loss function combined cross-entropy and KL Divergence terms:
+Self Distillation: We accelerated finetuning by using the original model’s predicted labels, akin to Knowledge Distillation techniques.
+The loss function combined cross-entropy and KL Divergence terms:
 $loss=\frac{1}{2}CrossEntropy(p,\hat{p})+\frac{1}{2}$ $KLDivergance(p,\tilde{p})$ (21)
@@
 After 5 epochs, the learning rate was reduced to $10^{-4}$.
-B.2 LayerCollapse vs. Knowledge Distillation
+C.3 LayerCollapse vs. Knowledge Distillation
 For both LayerCollapse and Knowledge Distillation, the target architectures were identical.
-The LayerCollapse setup followed the procedure outlined in Section B.1.
-The distillation process employed the SGD optimizer with an initial learning rate of $0.001$, reduced to $0.0001$ upon plateau.
-Experiments concluded when no further progress was achievable.
-B.3 Ablation Study of Regularization
+The LayerCollapse setup followed the procedure outlined in Section C.2.2.
+The distillation process employed the SGD optimizer with an initial learning rate of $0.001$, reduced to $0.0001$ upon plateau, and concluded when no further progress was possible.
+C.4 Study of Regularization on VGG
 To ensure comparability, all models, including those without LayerCollapse, shared the same training settings.
@@
 Hyper-parameter tuning was not a focus of this study.
-B.4 LayerCollapse in NLP
-We utilized the base Bert model from Hugging Face for NLP experiments. Finetuning employed the Adam optimizer with a learning rate of $0.00002$.